# DeepVoting: Learning Voting Rules with Tailored Embeddings

### v1

@@ 3 Preliminaries
Given a profile, the tournament embedding $T_{T}$ produces a $m\times m$ matrix $M$ where $M[j,k]=1$ if a majority of voters prefer $j\succ_{i}k$, $M[j,k]=0$ if a majority prefer $k\succ_{i}j$, and $M[j,k]=\frac{1}{2}$ if an equal number of voters prefer each candidate (when $n$ is even), for candidate pairs $j,k\in C$.
@@
Given a profile, the weighted tournament embedding $T_{WT}$ produces a $m\times m$ matrix $M$ where $M[j,k]=|\{i\in V:j\succ_{i}k\}|$ for $j,k\in C$.
@@
Given a profile, the rank frequency embedding $T_{RF}$ produces a $m\times m$ matrix $M$ showing how many voters rank each candidate $c\in C$ in each position $k\in[m]$ where $M[c,k]=|\{i\in V\text{ s.t. }\succ_{i}^{c}=k\}|$.
There is a tension in the literature between rules that use positional information, like scoring rules, and those that rely on majoritarian or pairwise comparison information, like tournament rules Brandt et al. (2014). Note $T_{RF}$ maintains positional information while $T_{WT}$ and $T_{T}$ maintain majoritarian.
Probabilistic Social Choice Functions
We now define our PSCFs. Where necessary, we always break ties lexicographically. Definitions for IRV and Black’s rule are in Appendix A. We include two rules that are typically classified as scoring rules in the voting literature, Borda and Plurality. The outcome of any scoring rule can be exactly computed from $T_{RF}$.
@@
The Borda score of candidate $c\in C$ from profile $X$ is $B(c)=\sum\limits_{i\in V}(m-x_{i}(c))$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}B(c)$ be the subset of candidates with maximum Borda score. The probabilistic Borda rule returns the lottery $U(W(X))$. [^1] Referred to as Borda Max by Endriss (2017).
Definition 5 (Plurality).
The rest of our rules are not scoring rules. Copeland is typically classified as tournament rule since it’s outcome can be computed directly from $T_{T}$ Brandt et al. (2014).
@@
Let $W$ be the subset of candidates whose maximum weight incoming edge is minimal in $G_{X}$. The probabilistic Simpson-Kramer work returns the lottery $U(W(X))$.
Definition 8 (Schulze).
@@
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality $\times$ $\times$
Borda ? $\times$
Copeland $\times$
Schulze $\times$ $\times$
Ranked Pairs $\times$ $\times$
Simpson-Kramer $\times$ $\times$
IRV $\times$ ? $\times$
Black’s Rule $\times$ ? $\times$
Table 1: PSFC preservation under embedding
@@ 4 PSCF Preservation Under Embedding
Embedding $T$ preserves PSCF $f$ if and only if for all pairs of profiles $X,\hat{X}\in\mathcal{X}$ we have $T(X)=T(\hat{X})\Rightarrow f(X)=f(\hat{X})$.
Some embeddings preserve strictly greater information than others. For example, lexicographically sorting the preference orders in a profile preserves all information necessary to compute $T_{WT}$, and $T_{WT}$ preserves all information necessary to compute $T_{T}$ from a profile. This implies that if $T_{T}$ preserves a function $f$, then $T_{WT}$ must preserve $f$ as well.
@@
Since we can compute $T_{T}$ from $T_{WT}$, $T_{T}$ can only preserve a PSCF if $T_{WT}$ does as well. However, the reverse does not hold. $T_{WT}$ may preserve PSCFs that are not preserved by $T_{T}$. If an embedding preserves a PSCF, then the PSCF is perfectly learnable from the embedding.
Table 1 shows which PSCFs are preserved by our embeddings. See Appendix B for proofs of the negative results in Table 1 where PSCFs are not preserved. The question marks in the table represent rule-embedding pairs for which we could not find such a counterexample from brute force search, and yet do not have a proof that preservation holds.
Figure 2: L1 rule loss on validation set of random profiles for Plurality, Borda, Copeland, and Schulze per epoch.
@@ 5 Learning Lotteries from PSCFs
In our first set of experiments, we show that with proper embeddings we can learn PSCFs that generalize common voting rules using network architectures with few parameters. We train each of our 21 rule-embedding pairs separately to compare their performance, and show how the choice of embedding must correspond to the choice of rule.
@@
We train our PSCFs on profiles with $n=29$ voters and $m=7$ candidates. For all experiments, profiles are generated uniformly at random, reflecting the impartial culture assumption Black et al. (1958). Like Firebanks-Quevedo (2020), we use the Whalrus package to implement our voting rules. [^2] https://pypi.org/project/whalrus/
Embeddings afford three key advantages: (1) They reduce the size of the input layer of our neural net, which is a fully connected layer, and therefore greatly reduce the number of model weights. All three of our embeddings compress the $n\times m$ profile to an $m\times m$ matrix representation, so the same MLP architecture can be used for all training runs. (2) When a rule is paired with an appropriate embedding, the embedding preserves all the information necessary to learn the rule and removes unnecessary information. (1) and (2) mean that we learn PSCFs faster and more accurately than previous work. (3) Input size no longer depends on the number of voters, which lends itself to better scaling. For $T_{RF}$ and $T_{WT}$ we normalize the input by dividing all elements in the embedding by $n$, e.g. the elements of the $T_{WT}$ now represent the fraction of voters who prefer one candidate to another $\frac{d(a,b)}{n}$ for $a,b\in C$.
We emulate the MLP architecture of Anil and Bao (2021), with 5 fully-connected layers of 120 nodes, ReLU activation functions, and a Softmax layer for the output. The key difference is that our network takes in embedded profiles so the size of our input layer is $m^{2}$ compared to their $nm^{2}$. This brings our total number of model parameters down to just under 50K vs. 200K. We train our models on a set of 48,000 randomly sampled profiles in batches of size 32 for 1000 epochs, for a total of 1.5M gradient steps. The use of embeddings also allows us to test our MLP model on larger profiles without increasing the size of the network, which Anil and Bao (2021) were unable to do for their MLP model, and only provided for their more sophisticated architectures. We trained each model on NVIDIA Volta V100 GPUs using PyTorch, with each run taking $\approx$1 hour. We used the Adam optimizer for each run with an initial learning rate of 0.001, tuning on plateau (patience = 10, factor = 0.5, min_lr = 0.01). We refer to the L1 distance between our model output and the PSCF lottery on a profile as the rule loss. Rule losses presented in Table 2 and Table 3 are from a test set of 10,000 random profiles sampled independently of the training data. All of these models were trained to minimize rule loss for their respective PSCFs.
@@
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.000706 0.154674 0.150783
Borda 0.002114 0.001468 0.065766
Copeland 0.070568 0.058405 0.000737
Schulze 0.071135 0.060596 0.044539
Simpson-Kramer 0.072062 0.056812 0.045097
IRV 0.092131 0.109611 0.079310
Black’s Rule 0.029726 0.029520 0.037891
Table 2: Average rule loss testing models trained to learn PSCFs on uniformly random profiles $(m=7,n=29)$.
Figure 2 shows plots of our validation losses during training, using data not used during training, to illustrate our observations. The plots for the remaining rules can be found in Appendix C.
The scoring rules, Plurality and Borda, learn very rapidly using $T_{RF}$ with rule losses converging to zero in a small number of epochs, as expected because $T_{RF}$ preserves all scoring rules. For plurality, we see a non-zero plateau for $T_{WT}$ and $T_{T}$, as they do not preserve all information needed to learn the rule, and so there is a non-zero lower bound to the error rate.
One of our most surprising results is that Borda appears to learn just as well from $T_{WT}$ as from $T_{RF}$. This raises the question of whether $T_{WT}$ preserves Borda. To our knowledge, there is no known explicit algorithm for computing the Borda outcome from a weighted tournament. We pose the same question for IRV and Black’s Rule with $T_{WT}$. In Social Choice there is thought to be a tension between positional information, such as that contained in $T_{RF}$, and pairwise comparisons. For example, no scoring rule can be Condorcet-consistent. So if Borda can be computed from weighted tournament graphs, as suggested by our learning experiments, this would be a surprising result. Our search for counterexamples with small numbers of voters and candidates did not yield any results, but this search was limited by computational intractability.
Challenge.
Does the weighted tournament embedding preserves Borda, IRV, or Black’s rule?
As expected, the Copeland rule learns rapidly with the $T_{T}$, converging to near zero loss quickly as Copeland can be computed from $T_{T}$. While any rule that can be computed from $T_{T}$ can also be computed from $T_{WT}$, what we see is that the Copeland rule loss converges to zero much more slowly for the $T_{WT}$, failing to reach the same loss as $T_{T}$ in our experiments after 1000 epochs. We make a similar observation for the Schulze rule. However, unlike Copeland, Schulze can be computed exactly from $T_{WT}$ but not $T_{T}$. This is why we see the loss with $T_{T}$ plateau at a nonzero value for Schulze in Figure 2. And yet, the rule loss for Copeland with $T_{WT}$ fails to surpass $T_{T}$ after 1000 epochs. This highlights both the benefits of choosing the right embedding for the rule, and that embeddings containing more information may lead to slower learning. Similar observation can be made for the Simpson-Kramer rule. For several of our rules, when using $T_{WT}$, the rule does not appear to finish learning after 1000 epochs. Ultimately, no single embedding performs best across all rules.
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.033898 0.173783 0.188456
Borda 0.016891 0.014817 0.081727
Copeland 0.085109 0.085285 0.000763
Schulze 0.089108 0.100582 0.149869
Simpson-Kramer 0.089382 0.100518 0.054845
IRV 0.114596 0.121245 0.094002
Black’s Rule 0.042577 0.040671 0.206753
Table 3: Rule loss on models trained on random profiles $(m=7,n=29)$ tested on profiles with $n=199$ voters.
Scaling With Number of Voters
Our embeddings give us the ability to work with profiles with different numbers of voters. Although we trained our models only with profiles with 29 voters, we can test with larger and smaller numbers of voters to check generalizability. We test first on profiles with 199 votes (Table 3), and then again with only 9 voters (Table 6 in Appendix D). We see that there is an increase in the loss in relative terms for all rules, but most losses remain quite low compared to their initial losses during training. We note that the Copeland rule with $T_{T}$ scaled the best, with low error in absolute terms across all treatments. All of the rules besides Plurality and Borda generalized better in relative terms with $T_{RF}$ than did Plurality and Borda, however Plurality and Borda maintained the lowest errors with that embedding. The rule-embedding pairs with the highest errors initially, also tended to scale quite well in relative terms. And all rules performed better with 9 voters than with 199 for all embeddings. When scaling to 199 voters, Schulze and Black’s rule did exceptionally poorly with $T_{T}$. Once again, this highlights the importance of choosing an embedding that fits the rule, corresponding to the learning objective. No single embedding does better than the others for scaling across the board.
Figure 3: Validation losses for models trained to learn PSCFs with $m=7,n=29$ and retrained to learn Participation with combined loss on profiles with $m=7,n=9$.
@@ 6 Resisting the No Show Paradox
PSCFs based on voting rules can be vulnerable to the No Show Paradox, where a voter prefers the outcome yielded by a rule when they do not vote. The voter therefore has an incentive to abstain rather than report their true preference. A rule for which this cannot occur is said to satisfy the Participation axiom. We now employ transfer learning, taking our already trained models from Section 5 and retraining them with a loss function that adds a term for Participation loss, which is our relaxation of the binary participation axiom to a continuous loss function.
For all definitions below, let $P_{X}$ be a probability distribution derived from profile $X$ by some PSCF $f$ (implicit), and let $P_{X}(c)$ be the probability assigned to candidate $c$. Where the specific profile is not relevant, we will denote simply by $P(c)$ the probability assigned to candidate $c$ by a lottery $P$. We now provide a loss function that measures how well a PSCF satisfies participation or how close a PSCF comes to satisfying participation. We use stochastic dominance to model a voter’s preference between two lotteries based on their preference order over candidates.
@@
We say that a voter’s abstention leads to an outcome ($P$) they prefer if the new outcome stochastically dominates the outcome ($Q$) that would derive from the true profile, with respect to the voter’s ordering of the candidates $\sigma=x_{i}$. We want our PSCF to be strategyproof with respect to the limited class of strategic abstentions.
@@
Since Participation is a binary condition for a PSCF, to learn PSCFs that resist the No Show Paradox, we define a non-binary loss function for Participation based on stochastic dominance.
@@
$T_{RF}$ $T_{WT}$ $T_{T}$
Plurality 0.282162 0.412889 0.411734
Borda 0.409378 0.408852 0.423503
Copeland 0.399701 0.380584 0.389514
Schulze 0.413933 0.388911 0.395796
Simpson-Kramer 0.403704 0.395422 0.393872
IRV 0.409915 0.413539 0.426163
Black’s Rule 0.416672 0.415060 0.409416
Table 4: Participation Loss of models trained to learn PSCFs on profiles of size ($m=7$, $n=29$), tested on 1000 random profiles of size ($m=7$, $n=9$).
The results in Table 4 are surprising. The single-winner versions of Borda, Plurality, and Simpson-Kramer do not suffer from the No Show Paradox, and yet our learned models for the PSCFs based on these rules show similar Participation losses to our other rules. Plurality showed a lower initial Participation loss when trained with $T_{RF}$, but not drastically lower. This may be due to our choice of the maximum stochastic dominance loss over all voters for a profile rather than, say, the sum of such losses across all voters. The rest of our rules are known to suffer from the paradox Pérez (2001), although the paradox does not arise frequently Brandt et al. (2019). And it is known that Condorcet-consistency is incompatible with Participation when there are at least 4 candidates and at least 12 voters Brandt et al. (2017). By our measure, almost all learned rules have similar initial Participation loss in Table 4.
@@
Retraining uses the same architectures and setup as the initial training. However, we retrain on 900 random profiles with 7 candidates and 9 votes as the No Show Paradox is more likely to occur with fewer voters. Our ability to change the numbers of voters, without padding the profile, is a benefit of our embeddings. To compute our loss, we add Participation Loss and the original L1 rule loss for each profile. We retrain for 200 total epochs per model.
Using fewer voters for training is also more computationally efficient, which is important because computing losses based on $n$ alternative profiles for each profile increases the runtime by $O(n)$. This is a major challenge for all inter-profile axioms that involve counterfactual comparisons. The added complexity hinges on how many different profiles must be considered to evaluate whether an axiom is satisfied Schmidtlein (2022); Schmidtlein and Endriss (2023)
We did not retrain all of our models for Table 5. Instead, for each rule, we retrained models using (1) the embedding(s) known to preserve the rule, if any, and (2) whichever embedding(s) performed best in our initial experiment. In Figure 3 we plot the epoch losses from an independent validation set with 100 disjoint random profiles. Each plot shows the Participation loss, rule loss (L1), and the sum of the two used as the combined loss function for training. Additional results are in Appendix E.
@@
The three rules we trained and retrained using $T_{RF}$ – Plurality, Borda, and Black’s Rule – all learned to minimize their Participation loss efficiently, with little or no increase in the rule loss, as shown in Table 5. Plurality learned most efficiently, being the only experiment in which the axiom loss dropped below the rule loss within 200 epochs. Despite Borda having higher Participation loss initially than we might have expected due to single-winner Borda satisfying Participation, it learned to evade the No Show paradox quite easily. Among the rules trained using $T_{T}$ and $T_{WT}$, Simpson-Kramer minimized the Participation loss most efficiently with $T_{T}$, which is not too surprising as single-winner Simpson-Kramer satisfies participation. However, the rule loss does not converge to zero at the same rate, even with $T_{WT}$. It is surprising that learning is so much poorer for Simpson-Kramer with $T_{WT}$ than $T_{T}$. The other rules retrained with $T_{T}$ and $T_{WT}$ struggled to learn to satisfy Participation without increasing rule loss within 200 epochs. This may be because appearances of the No Show Paradox are quite rare, occurring in about 4% of profiles Brandt et al. (2019).
Rule Embedding Participation Loss Rule Loss
Plurality $T_{RF}$ 0.078363 0.054809
Borda $T_{RF}$ 0.254161 0.029481
Borda $T_{WT}$ 0.239112 0.035640
Borda $T_{T}$ 0.432263 0.061900
Copeland $T_{T}$ 0.127699 0.039600
Schulze $T_{WT}$ 0.175672 0.089309
Schulze $T_{T}$ 0.065567 0.058519
Simpson-Kramer $T_{WT}$ 0.136363 0.105944
Simpson-Kramer $T_{T}$ 0.019293 0.062169
IRV $T_{WT}$ 0.165605 0.113929
IRV $T_{T}$ 0.229106 0.093015
Black’s Rule $T_{RF}$ 0.193693 0.039200
Black’s Rule $T_{WT}$ 0.215214 0.066647
Table 5: Participation Loss and Rule Loss (L1) of selected models retrained on profiles of size ($m=7$, $n=9$), averaged over 160 test profiles.

### v2

@@ 3.1 Embeddings
The tournament embedding $T_{T}$ yields a $m\times m$ matrix $M$ where $M[j,k]=1$ if a majority of voters prefer $j\succ_{i}k$, $M[j,k]=0$ if a majority prefer $k\succ_{i}j$, and $M[j,k]=\frac{1}{2}$ if an equal number of voters prefer each candidate (when $n$ is even), for candidate pairs $j,k\in C$.
@@
The weighted tournament embedding $T_{WT}$ yields a $m\times m$ matrix $M$ where $M[j,k]=|\{i\in V:j\succ_{i}k\}|$ for $j,k\in C$.
@@
The rank frequency embedding $T_{RF}$ yields a $m\times m$ matrix $M$ of how many voters rank each candidate $c\in C$ in each position $k\in[m]$ where $M[c,k]=|\{i\in V\text{ s.t. }\succ_{i}^{c}=k\}|$.
There is a tension in the literature between rules that use positional information, like scoring rules, and those that rely on majoritarian or pairwise comparison information, like tournament rules (Brandt, Brill, and Harrenstein 2014). Note $T_{RF}$ maintains positional information while $T_{WT}$ and $T_{T}$ are majoritarian. At times we also refer to a concatenation of all three embeddings which we refer to as $T_{CO}$.
Definition 4 (Combined Embedding).
The combined embedding $T_{CO}$ is the $3m^{2}$ concatenation of $T_{RF},T_{T},T_{WT}]$.
3.2 Probabilistic Social Choice Functions
We now define our PSCFs. Where necessary, we always break ties lexicographically. Definitions for Plurality, Schulze, Instant Runoff Voting (IRV) and Black’s rule can be found in the Appendix. Two rules that are typically classified as scoring rules, Borda and Plurality. The outcome of any scoring rule can be exactly computed from $T_{RF}$.
@@ 3.2 Probabilistic Social Choice Functions
The Borda score of candidate $c\in C$ from profile $X$ is $B(c)=\sum\limits_{i\in V}(m-x_{i}(c))$. Let $W(X)=\operatorname*{arg\,max}\limits_{c\in C}B(c)$ be the subset of candidates with maximum Borda score. The probabilistic Borda rule returns the lottery $U(W(X))$ (Referred to as Borda Max by Endriss (2017)).
The rest of our rules are not scoring rules. Copeland is a tournament rule since it’s outcome can be computed directly from $T_{T}$ (Brandt, Brill, and Harrenstein 2014).
@@
Let $W$ be the subset of candidates whose maximum weight incoming edge is minimal in $G_{X}$. The probabilistic Simpson-Kramer rule returns the lottery $U(W(X))$.
@@ 4 PSCF Preservation Under Embedding
Embedding $T$ preserves PSCF $f$ if and only if $f(X^)T(X)=T(\hat{X})\Rightarrow$ $f(\hat{X})$ for all $X,\hat{X}\in\mathcal{X}$.
Some embeddings preserve strictly more information than others. For instance, $T_{WT}$ preserves all information necessary to compute $T_{T}$ from a profile. This implies that if $T_{T}$ preserves a function $f$, then $T_{WT}$ must preserve $f$ as well.
@@
As we can compute $T_{T}$ from $_{WT}, TT$$_{T}$ can only preserve a PSCF if $T_{WT}$ does as well, the reverse does not hold. $T_{WT}$ may preserve PSCFs that are not preserved by $T_{T}$. If an embedding preserves a PSCF, then the PSCF is perfectly learnable from the embedding. Table 1 (green highlights) shows which PSCFs are preserved by these embeddings. See Appendix B for proofs of the negative results where PSCFs are not preserved. Each of these proofs consists of a counterexample with two profiles whose outcomes differ under the PSCF yet have the same embedding. Finally, many of our rules fall into Fishburn’s categorization: the winner of C1 rules (Copeland) can be computed using the information encoded in $T_{T}$, while the winner of C2 rules (Borda, Schulze, Simpson-Kramer, Black’s) can be computed from $T_{WT}$. A separate categorization of voting rules, positional scoring rules (Plurality, Borda), can be computed using only $T_{RF}$  (Brandt et al. 2016). IRV has recently been shown not to belong to either C1 or C2, and is not a positional scoring rule (Halpern, Hossain, and Tucker-Foltz 2024).
Figure 2: Validation Loss per epoch for each rule and embedding pair.
@@ 5 Learning Lotteries from PSCFs
First, we show that with suitable embedding we can learn PSCFs that generalize common voting rules using network architectures with few parameters. We train rule-embedding pairs separately for 32 combination of rule and embedding (3 embeddings and their concatenation) to compare their performance, and explore the rule-embedding tradeoff. Mohsin et al. (2022) use some of the same embeddings with XGBoost. However, we are the first to use them with MLPs, and hence they must be validated. MLPs also allow us to fine-tune these rules later in Section Section 6, which is not possible with XGBoost.
@@
We train our PSCFs on profiles with $n=44$ voters and $m=11$ candidates. For all experiments, profiles are sampled from the impartial culture distribution – i.e., rankings are generated uniformly at random (Black 1958). Like Firebanks-Quevedo (2020), we use the Whalrus package to implement our voting rules. [^2] https://pypi.org/project/whalrus/
Embeddings afford three key advantages: (1) They reduce the size of the input layer of our network, which is fully connected, and therefore greatly reduce the number of model weights. All three embeddings compress the $n\times m$ profile to an $m\times m$ matrix representation, so the same MLP architecture can be used for all training runs. (2) When a rule is paired with an appropriate embedding, the embedding preserves all the information necessary to learn the rule and removes unnecessary information. (1) and (2) mean that we learn PSCFs faster and more accurately than previous work. (3) Input size no longer depends on the number of voters, which lends itself to better scaling. For $T_{RF}$ and $T_{WT}$ we normalize by dividing all elements in the embedding by $n$, e.g., the elements of $T_{WT}$ represent the fraction of voters who prefer one candidate to another $\frac{d(a,b)}{n}$ for $a,b\in C$.
We emulate the MLP architecture of Anil and Bao (2021), with 5 fully-connected layers, the first hidden layer with 200 nodes, then four with 120 nodes, TanH activation functions, and a Softmax layer for the output. [^3] Discussion of other setups is in Appendix C. The key difference is that our network takes in embedded profiles so the size of our input layer is $m^{2}$ compared to their $nm^{2}$. This brings our total number of model parameters down to $\approx$100K vs. millions. We train our models on a set of 100,000 randomly sampled profiles in batches of size 32 for 1000 epochs, for a total of 1.5M gradient steps. The use of embeddings also allows us to test our MLP model on larger voter profiles without increasing the size of the network, which Anil and Bao (2021) were unable to do for their MLP model. We trained each model on NVIDIA A100 GPUs using PyTorch, with each run taking $\approx$4 hours. We used the Adam optimizer for each run with an initial learning rate of 0.001, tuning on plateau (patience = 50, factor = 0.5, min_lr = $1e^{-5}$).
We refer to the L1 distance between our model output and the PSCF lottery on a profile as the rule loss. Rule losses presented in Table 1 are from a test set of 10,000 random profiles sampled independently of the training data. All models are trained to minimize rule loss for their PSCFs.
@@
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.0 0.129 0.136 0.001
Borda 0.165 0.048 0.018 0.042
Copeland 0.088 0.0 0.163 0.0
Schulze 0.071 0.043 0.067 0.042
Simpson-Kramer 0.071 0.043 0.062 0.043
IRV 0.159 0.105 0.159 0.161
Black’s Rule 0.046 0.044 0.165 0.037
Ranked Pairs 0.161 0.051 0.161 0.07
Table 1: L1 Loss for each embedding on test data sampled from the Impartial Culture for models targeting each rule using $(m=11,n=44)$. Shaded cells indicate the embedding contains sufficient information to learn the rule perfectly.
Table 1 gives final validation set results and Figure 2 shows plots of our validation losses during training, using 10,000 samples from a held-out validation set, to demonstrate the effectiveness of learning for each rule-embedding pair. Plurality learns rapidly using $T_{RF}$ and $T_{CO}$ with rule losses converging to zero quickly. Our other positional scoring rule, Borda, has significant trouble learning from $T_{RF}$ despite the embedding having enough information to compute the Borda winner. For Plurality, we see some learning from $T_{WT}$ and $T_{T}$, but learning quickly plateaus as the embeddings do not preserve all information needed to learn the rule, and so there is a non-zero lower bound to the error rate. The Copeland rule learns rapidly with the $T_{T}$ and $T_{CO}$, converging to near zero loss quickly since it is preserved. While any rule that can be computed from $T_{T}$ can also be computed from $T_{WT}$, what we see is that the Copeland rule loss falls far more slowly with $T_{WT}$, failing to reach the same loss as $T_{T}$ in our experiments after 1000 epochs. We make a similar observation for the Schulze rule. However, unlike Copeland, Schulze can be computed exactly from $T_{WT}$ but not $T_{T}$. This is why we see the loss with $T_{T}$ plateau at a nonzero value for Schulze. Looking at Table 1 we can see the variation of final loss across all embeddings. While $T_{CO}$ contains all the information, in some cases we are able to more effectively learn from smaller embeddings. This highlights the challenges of working with neural networks, the benefits of choosing the right embedding for the rule, and that embeddings containing more information do not always help learning.
5.1 Comparison to Single-Winner
We now test our PSCFs for their accuracy in identifying the unique winners of each rule (when they exist) to directly compare with the four rules (Plurality, Borda, Copeland, and Simpson-Kramer) of Anil and Bao (2021). For each rule, we sample profiles with unique winners, obviating the problems of tie-breaking, and select the candidate with the highest probability mass as the winner. [^4] This rejection sampling method of Anil and Bao (2021) eliminates profiles with multiple winners, which may introduce artifacts into the accuracy measures. Our PSCFs do not share this problem. Note that our MLP architecture is the same as Anil and Bao (2021), only differing with better embeddings, and results are given in Table 2. We can see that our models learn Plurality, Borda, Copeland, and Simpson-Kramer extremely well. Our performance for these four rules is on par (Plurality, Borda) or better (Copeland, Simpson-Kramer) with the results of Anil and Bao (2021), and in some cases we are able to outperform even their more complex PIN architectures. For example, we learn Copeland perfectly (1.0) whereas across their four architectures their best performing model is 0.83; Simpson-Kramer our models (0.913) strictly outperform all of theirs (best of 0.80). Hence by leveraging embeddings we are able to learn rules as well as or better with models orders of magnitude smaller.
Target Rule $_{RF} TT$$_{T}$ $T_{WT}$ $T_{CO}$
Plurality 0.999 0.355 0.289 0.997
Borda 0.086 0.826 0.934 0.844
Copeland 0.552 1.0 0.087 1.0
Schulze 0.719 0.903 0.735 0.914
Simpson-Kramer 0.72 0.902 0.79 0.913
IRV 0.118 0.446 0.118 0.106
Black’s Rule 0.767 0.866 0.088 0.88
Ranked Pairs 0.102 0.814 0.102 0.67
Table 2: Accuracy of our models on each embedding using test data sampled from the Impartial Culture using $(m=11,n=44)$. Most accurate embeddings are shown in bold.
5.2 Beyond Impartial Culture
We now consider the generality of our training distribution against other preference distributions. As the Impartial Culture (IC) provides orders where candidates are ranked uniformly at random, all possible profiles have a non-zero probability of occurring. That is, given sufficient data, IC will generate profiles that could have been generated by all other distributions. However, recent (Boehmer et al. 2024, 2022; Szufa et al. 2025) and older (Mattei and Walsh 2013, 2017) work in the COMSOC community has illustrated the need to test on a wide variety of synthetic and real world preference distributions to ensure generalization. We test the empirical merit of this fact by evaluating our networks trained on IC preferences on test sets sampled from a wide range of distributions, including real-world preference data, complete results and definitions of distributions are in Appendix E.
Target Rule IC IAC Urn Mall. SP PrefL
Plurality .001 .001 .018 .012 .031 .014
Borda .028 .03 .077 .093 .135 .109
Copeland .0 .0 .001 .001 .0 .002
Schulze .017 .016 .029 .031 .018 .042
SK .017 .016 .028 .027 .02 .039
IRV .163 .162 .163 .164 .182 .162
Black’s .022 .023 .066 .085 .018 .103
RP .06 .06 .061 .06 .001 .063
Table 3: L1 Loss across preference distributions of networks trained using Impartial Culture and the $T_{CO}$ embedding.
Table 3 shows the loss of networks trained on $T_{CO}$ tested on 10,000 profiles from each other distribution. We include this evaluation for networks trained on other embeddings in Appendix E. In all cases, rules trained on $T_{CO}$ are able to generalize very effectively to new distributions. On highly structured preferences, such as the Single-Peaked distribution, some rules (e.g., Black’s, Ranked Pairs) have lower loss than on Impartial Culture preferences.
@@ 6 Resisting the No Show Paradox
Figure 3: Validation losses for models trained to learn PSCFs using $T_{CO}$ with $m=11,n=44$ and retrained to learn Participation. Combined loss shows the sum of rule and Participation Loss. Most rules converge quickly.
PSCFs based on voting rules can be vulnerable to the No Show Paradox, where a voter prefers the outcome yielded by a rule when they do not vote, giving an incentive to abstain. A rule for which this cannot occur is said to satisfy the Participation axiom. We now employ transfer learning, taking our trained models from Section 5 and retraining them with a loss function that adds a term for Participation loss.
For all definitions below, let $P_{X}$ be a probability distribution derived from profile $X$ by some PSCF $f$ (implicit), and let $P_{X}(c)$ be the probability assigned to candidate $c$. Where the specific profile is not relevant, we will denote simply by $P(c)$ the probability assigned to candidate $c$ by a lottery $P$. We use stochastic dominance to model a voter’s preference between two lotteries based on their preference order over candidates to define Participation for PSCFs.
@@
We say that a voter’s abstention leads to an outcome ($P$) they prefer if the new outcome stochastically dominates the outcome ($Q$) that would derive from the true profile, with respect to the voter’s ordering of the candidates $\sigma=x_{i}$. We want our PSCF immune to strategic abstentions.
@@
Since Participation is a binary condition for a PSCF, to learn PSCFs that resist the No Show Paradox, we define a non-binary loss function based on stochastic dominance.
@@
Fine-tuning uses the same architectures and setup as the initial training. However, we retrain on 1056 random profiles with 11 candidates and 44 voters as, like most manipulations, the No Show Paradox is more likely to occur with fewer voters (Xia and Conitzer 2008). Changing the numbers of voters, without padding the profile, is a benefit of our embeddings. We add Participation Loss and the original rule loss for each profile and retrain for 40 total epochs taking about 12 hours each. Using fewer voters for training is also more computationally efficient, which is important as computing losses based on $n$ alternative profiles for each profile increases the runtime by $O(n)$. This is a major challenge for all inter-profile axioms that involve counterfactual comparisons as it determines how many different profiles must be considered determine if an axiom is satisfied (Schmidtlein 2022; Schmidtlein and Endriss 2023).
Before FT After FT
Target Rule Rule Part. Rule Part.
Plurality 0.0005 0.0713 0.0294 0.0000
Borda 0.0381 0.0000 0.0432 0.0000
Copeland 0.0000 0.2220 0.0051 0.0000
Schulze 0.0304 0.2152 0.0265 0.0000
Simpson-Kramer 0.0330 0.0015 0.0323 0.0000
IRV 0.1629 0.0000 0.1629 0.0000
Black’s Rule 0.0309 0.0000 0.0488 0.0000
Ranked Pairs 0.0699 0.0000 0.0797 0.0000
Table 4: Loss of our $T_{CO}$ models before and after fine-tuning with Participation Loss using $(m=11,n=44)$.
@@
Figure 3 gives the training loss per epoch while Table 4 gives both the Rule Loss and Participation Loss of rules before and after fine-tuning evaluated on a disjoint 1056 profiles from the training set; additional results for all embeddings are in Appendix F. These results are interesting in several ways. First, the single-winner versions of Borda, Plurality, and Simpson-Kramer resist the No Show Paradox, but only our learned Borda PSCF is resistant, with others showing some loss before fine-tuning. The rest of our rules are known to suffer from the paradox (Pérez 2001), and it is known that Condorcet-consistency is incompatible with Participation when there are at least 4 candidates and 12 voters (Brandt, Geist, and Peters 2017), although the paradox does not arise frequently, only in about 4% of profiles (Brandt, Hofbauer, and Strobel 2019). This leads to one of the most interesting results, we see both Copeland and Schulze, Condorcet Consistent rules, able to be fine-tuned in a way that mostly preserves the rule loss, but is also (empirically) immune to the no show paradox. While these are only small scale test, they point an intriguing way forward for future research.